If a ball is thrown up with an initial velocity of 48ft per second. Its height s after t is given by the equation s = 48t-16ft^2. Find the maximum height attained by the ball and the time it takes for the ball to return to Earth
Answer by amarjeeth123(569)

s=48t-16t^2
Velocity v=ds/dt=48-32t
For maximum height v=0
48-32t=0
32t=48
t=48/32=1.5 sec
Maximum height s(max)=48(1.5)-16(1.5^2)=72-36=36 feet.
Total time to come back to Earth=1.5(2)=3 sec